# Entanglement Entropy of Free Fermions in Timelike Slices

### v3

@@ III.1 Zero temperature
By Eq. 13, This gives
@@
Note that $D_{mn}$ can be viewed as the spatial correlation matrix of a “lattice fermion model” with a “lattice constant” $\tau_{0}$, where $E$ is the “quasi-momentum” in the “Brillouin zone” $[-\frac{\pi}{\tau_{0}},\frac{\pi}{\tau_{0}}]$, and $n_{F}(E)=\Theta(E_{F}-E)$ for some Fermi energy $E_{F}$. This is equivalent to a “zero temperature Fermi sea” state occupying “quasi-momentum” $-\frac{E_{0}}{2}\leq E\leq E_{F}$. By the Calabrese-Cardy formula [12, 13, 14], the zero-temperature entanglement entropy of 1D gapless states in a spatial subregion of length $\ell$ is $S_{A}=\frac{c_{L}+c_{R}}{6}\ln(\ell/\ell_{c})$ with some constant $\ell_{c}$, where $c_{L}$ and $c_{R}$ are the left and right central charges, respectively. Therefore, the zero-temperature time-direction entanglement entropy $S_{A}$ at $\tau=\tau_{0}$ should resemble that of 1D gapless lattice fermion (nonchiral with $c_{L}=c_{R}=1$) in a length $t$ subregion:
@@
where $t_{c}$ is some constant.
@@
(iii) When $\tau\ll\tau_{0}$, or $K\gg t/\tau_{0}$, if we enlarge the range of energy $E$ from $[-\frac{E_{0}}{2},\frac{E_{0}}{2}]$ to an ancillary range $[-\frac{\pi}{\tau},\frac{\pi}{\tau}]$, we can think of $E$ as the “quasi-momentum” in the Brillouin zone $[-\frac{\pi}{\tau},\frac{\pi}{\tau}]$ of an “ancillary lattice fermion model” with lattice constant $\tau$, which we assume has an ancillary dispersion $\omega_{a}(E)$ in Fig. 3d. Assume the ancilla model has a “Fermi level” $\omega_{a}^{F}=\omega_{a}(E_{F})$. Meanwhile, the physical energy range $[-\frac{E_{0}}{2},\frac{E_{0}}{2}]$ plays the role of a “quasi-momentum” cutoff near the right “Fermi surface” in Fig. 3d, keeping only the right-moving chiral fermions. As a result, the correlation matrix $D$ resembles that of one right-moving chiral fermion mode. Thus, $S_{A}$ stabilizes into the Calabrese-Cardy formula with $c_{L}=0,c_{R}=1$:
$S_{A}\approx\frac{1}{6}\ln\left(\frac{t}{t_{c}^{\prime}}\right)\approx\frac{1}{6}\ln\left(\frac{K\tau}{t_{c}^{\prime}}\right).\quad(\tau\ll\tau_{0},T=0)$ (42)
III.2 Finite temperature
@@
At temperature $T>0$, as shown in Fig. 3c, $S_{A}$ for fixed total time $t$ still shows a triangle or trapezoid like curve with respect to $K$ when $K<t/\tau_{0}$ ($\tau>\tau_{0}$), and stabilizes into a linear relation when $K>t/\tau_{0}$ ($\tau<\tau_{0}$). The small $K$ slope obeys Eq. 37 with a similar reasoning.
@@
To characterize the entanglement between slice $A$ and its complement, we calculate the mutual information $\mathcal{I}$ in Eq. 16. For fixed total time $t$, $\mathcal{I}$ at temperature $T>0$ is shown in Fig. 3e, which shows stabilization at $K>t/\tau_{0}$ ($\tau<\tau_{0}$), similar to $S_{A}$ at $T=0$ (Fig. 3a). At $T=\infty$, one can show $\mathcal{I}\equiv 0$. For fixed $\tau<\tau_{0}$, we find $\mathcal{I}$ increases at small $t$, and saturates as $t\gtrsim\beta=T^{-1}$, as shown in Fig. 3f.
III.3 Non-eigenstates
@@
We now study $S_{A}$ in time direction slice $A$ of Fig. 1b at site $m=0$ for pure non-eigenstates. We first consider state
@@
which has $N_{f}$ consecutive occupied sites. As shown in Figs. 4c and 4d with $N_{f}=100$, the behavior of $S_{A}$ changes sharply as $t$ reaches two times $t_{L}=\frac{N_{f}a_{0}}{3v_{\text{max}}}\approx 67$ and $t_{R}=\frac{2N_{f}a_{0}}{3v_{\text{max}}}\approx 133$, which correspond to the times for signals at two edges of the occupied interval to reach slice $A$. For $t<\min(t_{L},t_{R})$, $S_{A}$ versus $K$ with fixed $t$ (Fig. 4c) or fixed $\tau$ (Fig. 4d) remains almost zero for $t<\min(t_{L},t_{R})$, resembling that of a fully occupied state. When $t>\text{min}(t_{L},t_{R})$, Slice $A$ sees the state as partially filled, and $S_{A}$ starts to grow.
@@ V DISCUSSION
In this paper, we have defined a generalization of entanglement entropy in a generic spacetime slice for free fermion lattice models, as given by Eq. 15. We note that this spacetime entanglement entropy we defined easily applies to free bosons [52], which can similarly be calculated from two-point correlation functions.
For 1D free lattice fermion tight-binding model, we find the entanglement entropy in a time-direction slice approaches Eq. 42, which holds when $\tau<\tau_{0}$. This indicates that the zero temperature time-direction entanglement entropy $S_{A}$ of 1D free lattice fermions resembles the chiral Calabrese-Cardy formula $S_{A}=\frac{1}{6}\ln\left(\frac{t}{t_{c}^{\prime}}\right)$ in the continuous time limit $\tau\rightarrow 0$, given that the system size $L\gg K=\frac{t}{\tau}+1$. A continuous time limit also exists for mutual information $\mathcal{I}$ at finite temperature. When $K\geq L$, the Hilbert space $h_{A}$ will saturates to $h_{\text{tot}}$ (unless constrained by symmetries), leading to $S_{A}=S_{\text{tot}}$ [51].
@@ I Fermion Operator Orthonormalization
We give the details of orthonormalizing the free fermion creation and annihilation operators in an arbitrary spacetime slice $A$ in this section. Consider the full system with a total number of fermion modes $L$. Given $K$ single-particle annihilation operators $c_{\mathbf{r}_{1}}(t_{1}),c_{\mathbf{r}_{2}}(t_{2}),\cdots c_{\mathbf{r}_{K}}(t_{K})$ in slice $A$, we can express them as

### v4

@@ III.2.2 The $\tau=\tau_{0}$ case: duality to the spatial case
namely, $B=I$ is simply the identity matrix. This indicates that the fermion operators $c_{0}(t_{n})$ at different time $t_{n}$ already form an orthogonal basis, and thus we can choose the matrix $M=I$. By Eq. 15, this gives
@@
Note that this $D_{mn}$ can be viewed as the spatial correlation matrix of a “lattice fermion model” with a “lattice constant” $\tau_{0}$, where $E$ is the “quasi-momentum” in the “Brillouin zone” $[-\frac{\pi}{\tau_{0}},\frac{\pi}{\tau_{0}}]$, and $n_{F}(E)=\Theta(E_{F}-E)$ for some Fermi energy $E_{F}$. Therefore, the time-direction correlation matrix $D_{mn}$ is equivalent to the spatial correlation matrix of a “zero temperature Fermi sea” state of a 1D free fermion lattice model occupying the interval of “quasi-momentum” $-\frac{E_{0}}{2}\leq E\leq E_{F}$. In other words, we see that the time-direction entanglement entropy $S_{A}$ at $\tau=\tau_{0}$, which is a function of correlation matrix $D$, is dual to the conventional spatial entanglement entropy of a free fermion lattice model.
By the Calabrese-Cardy formula [12, 13, 14], the zero-temperature entanglement entropy of gapless states of infinite 1D space in a spatial subregion of length $\ell$ is $S_{A}=\frac{c_{L}+c_{R}}{6}\ln(\ell/\ell_{c})$ with some ultraviolet (UV) cutoff constant $\ell_{c}$, where $c_{L}$ and $c_{R}$ are the left and right central charges, respectively. Therefore, the above duality indicates that our zero-temperature time-direction entanglement entropy $S_{A}$ at $\tau=\tau_{0}$ should resemble that of the 1D gapless lattice fermion (which is nonchiral with $c_{L}=c_{R}=1$) in a length $t$ subregion (in the limit $L\rightarrow\infty$):
@@
where $t_{c}$ is some UV cutoff constant. This agrees with our observations in Fig. 3c. The similarity between time-direction entanglement entropy in Eq. 52 and the conventional spatial entanglement entropy of free lattice fermions also suggests that this $\tau=\tau_{0}$ case here might be similar to the dual unitary circuits [56, 57, 58].
III.2.3 The $\tau\ll\tau_{0}$ case: the continuum time limit
When $\tau\ll\tau_{0}$, or $K\gg t/\tau_{0}$, we are approaching the continuum time limit where the neighboring points are close in time.
To explain the numerical results in Fig. 3d, we imagine the discrete time points at site 00 can be understood as an “ancillary lattice fermion model” with lattice constant $\tau$, and the time direction is playing the role of “spatial direction” of this “ancillary model”. In this picture, the original energy $E$ can be viewed as the “quasi-momentum” of this “ancillary model”, which lives in the Brillouin zone of this “ancillary model” which is an interval $[-\frac{\pi}{\tau},\frac{\pi}{\tau}]$. Assume this “ancillary model” has an ancillary dispersion $\omega_{a}(E)$ as shown in Fig. 4a, and a “Fermi level” $\omega_{a}^{F}=\omega_{a}(E_{F})$, where $E_{F}$ is the physical Fermi energy, or the “Fermi momentum” of this “ancillary model”. Then the physical Fermi sea state, which has levels $E<E_{0}$ fully occupied, maps to the “Fermi sea state” of the “ancillary model” with levels $\omega_{a}(E)<\omega_{a}^{F}$ fully occupied.
@@ III.2.3 The $\tau\ll\tau_{0}$ case: the continuum time limit
Meanwhile, physically, the energy $E$ can only take values in the energy range of the physical band $[-\frac{E_{0}}{2},\frac{E_{0}}{2}]$, which is much smaller than the Brillouin zone of the “ancillary model”. Therefore, the physical energy range $[-\frac{E_{0}}{2},\frac{E_{0}}{2}]$ plays the role of a “quasi-momentum” cutoff near the right “Fermi surface”, as shown by the solid line in Fig. 4a, keeping only the right-moving chiral fermions of the “ancillary model”. Thus, the correlation matrix $D$ resembles that of one right-moving chiral fermion mode, and accordingly, we expect the time-direction entanglement entropy $S_{A}$ in the $\tau\ll\tau_{0}$ limit to stabilize into the Calabrese-Cardy formula for a chiral fermion mode with $c_{L}=0,c_{R}=1$ (in the limit $L\rightarrow\infty$), namely:
$S_{A}\approx\frac{1}{6}\ln\left(\frac{t}{t_{c}}\right)\approx\frac{1}{6}\ln\left(\frac{K\tau}{t_{c}}\right).\quad(\tau\ll\tau_{0},T=0)$ (53)
where again $t_{c}$ is some UV cutoff constant.
More intuitively, the result of Eq. 53 in the continuum time limit can be understood from the picture of Fig. 4b-c. In an infinite 1D lattice model, the ground state, when viewed in the momentum $k$ space, is a Fermi sea state occupying the momentum interval $[-k_{F},k_{F}]$ (Fig. 4b), which has two Fermi points at $k_{F}$ and $-k_{F}$. Accordingly, the spatial entanglement entropy is $S_{A^{\prime}}=\frac{1}{3}\ln(\ell/\ell_{c})$ for a spatial subregion $A^{\prime}$ of length $\ell$, where the coefficient $\frac{1}{3}$ is given by $\frac{1}{6}$ times the total number of Fermi points in the momentum space. This is because each Fermi point contributes a chiral fermion degree of freedom, contributing chiral central charge $1$. In analogy, to understand the time-direction entanglement entropy, we should view the ground state in the energy $E$ space, which is a Fermi sea state occupying the energy range $E<E_{F}$ (Fig. 4c). In this case, there is only one Fermi point $E_{F}$ in the energy space (the lower bound $-\frac{E_{0}}{2}$ plays the role of the UV cutoff rather than a Fermi point, as explained by Fig. 4a). Accordingly, the time-direction entanglement entropy $S_{A}=\frac{1}{6}\ln(t/t_{c})$, and the coefficient $\frac{1}{6}$ is given by $\frac{1}{6}$ times the number of Fermi points in the energy space.
Therefore, we see that the time-direction entanglement entropy in the continuum time limit is probing the “effective central charge” in the time direction, which is the number of Fermi points in the energy space. This piece of information is different from that given by the spatial entanglement entropy.
Refer to caption(a)
(b)Refer to caption(c)Refer to caption
Figure 4: (a) The ancillary dispersion for explaining $S_{A}$ when $\tau<\tau_{0}$ (“anc” for ancillary, “emp” for empty). (b) For 1D gapless lattice fermion, the zero temperature Fermi sea state has two Fermi points at ±$\pm k_{F}$ in the momentum space. Accordingly, the Calabrese-Cardy formula for the spatial entanglement entropy is $S_{A^{\prime}}=\frac{1}{3}\ln\left(\frac{l}{l_{c}}\right)$. (c) In the energy space, the zero temperature state has single-particle states with $E\leq E_{F}$ are filled, so there is only one Fermi point $E_{F}$ in the energy domain. Accordingly, the time-direction entanglement entropy in the continuum time limit is $S_{A}=\frac{1}{6}\ln\left(\frac{t}{t_{c}}\right)$.
III.3 Finite temperature
@@
At temperature $T>0$, the total state is in the thermal mixed state. As shown in Fig. 3c, $S_{A}$ for fixed total time $t$ still shows a triangle or trapezoid like curve with respect to $K$ when $K<t/\tau_{0}$ ($\tau>\tau_{0}$), and stabilizes into a linear relation when $K>t/\tau_{0}$ ($\tau<\tau_{0}$). This shows that the time-direction entanglement entropy $S_{A}$ shows a volume law in $t$ (namely, linear in $t$) at finite temperature. The small $K$ (namely, $\tau\gg\tau_{0}$ for fixed total time $t$) slope obeys Eq. 47 with a similar reasoning.
@@
We also calculate the mutual information $\mathcal{I}$ in Eq. 18. For fixed total time $t$, $\mathcal{I}$ at temperature $T>0$ is shown in Fig. 3e, which shows stabilization at $K>t/\tau_{0}$ ($\tau<\tau_{0}$), similar to $S_{A}$ at $T=0$ (Fig. 3a). At $T=\infty$, it is straightforward to show that $\mathcal{I}\equiv 0$. For fixed $\tau<\tau_{0}$, we find $\mathcal{I}$ increases at small $t$, and saturates as $t\gtrsim\beta=T^{-1}$, as shown in Fig. 3f.
III.4 Non-eigenstates
@@
We now study the time-direction entanglement entropy $S_{A}$ in time direction slice $A$ of Fig. 1b at site $m=0$ for pure non-eigenstates. We first consider state
@@
which has $N_{f}$ consecutive occupied sites. As shown in Figs. 5c and 5d with $N_{f}=100$, the behavior of $S_{A}$ changes sharply as $t$ reaches two times $t_{L}=\frac{N_{f}a_{0}}{3v_{\text{max}}}\approx 67$ and $t_{R}=\frac{2N_{f}a_{0}}{3v_{\text{max}}}\approx 133$, which correspond to the times for signals at two edges of the occupied interval to reach slice $A$.
The above behaviors of $S_{A}$ can be understood as follows. For $t<\min(t_{L},t_{R})$, $S_{A}$ versus $K$ with fixed $t$ (Fig. 5c) or fixed $\tau$ (Fig. 5d) remains almost zero for $t<\min(t_{L},t_{R})$, resembling that of a fully occupied state. This is because the signals from the edges of the occupied interval has not yet propagated to the time direction slice $A$, and the slice $A$ effectively sees a fully occupied state with zero entropy. When $t>\text{min}(t_{L},t_{R})$, the time direction slice $A$ starts to see the state as partially filled, and $S_{A}$ starts to grow.
@@ V DISCUSSION
In this paper, we have defined a generalization of entanglement entropy in a generic spacetime slice for free fermion lattice models, as given by Eq. 17. We also discussed calculating this spacetime entanglement entropy for free fermion lattice models with pairings. A future question is to generalize this spacetime entanglement entropy we defined to free boson systems, which should also be in principle calculable from two-point correlation functions due to the Wick’s theorem [52].
For 1D free lattice fermion tight-binding model, we find the zero temperature time-direction entanglement entropy in a time-direction slice approaches $S_{A}=\frac{1}{6}\ln\left(t/t_{c}\right)$ (Eq. 53) when $\tau<\tau_{0}$. This indicates that the zero temperature time-direction entanglement entropy $S_{A}$ of 1D free lattice fermions in the continuous time limit $\tau\rightarrow 0$ resembles the spatial entanglement entropy of a chiral fermion mode, given that the system size $L\gg K=\frac{t}{\tau}+1$. We showed that the coefficient $\frac{1}{6}$ is probing the number of Fermi points in the energy space. A continuous time limit also exists for mutual information $\mathcal{I}$ at finite temperature. Moreover, in the special case $\tau=\tau_{0}$, we find the time-direction entanglement entropy is given by $S_{A}=\frac{1}{3}\ln\left(t/t_{c}\right)$, which is dual to the spatial entanglement entropy. Lastly, we note that when $K\geq L$, the Hilbert space $h_{A}$ will saturates to $h_{\text{tot}}$ (unless constrained by symmetries), leading to $S_{A}=S_{\text{tot}}$ [51].
@@ I Fermion Operator Orthonormalization
We give the details of orthonormalizing the free fermion creation and annihilation operators in an arbitrary spacetime slice $A$ in this section.
I.1 The case with fermion number conservation
Consider the full system with a total number of fermion modes $L$. Given $K$ single-particle annihilation operators $c_{\mathbf{r}_{1}}(t_{1}),c_{\mathbf{r}_{2}}(t_{2}),\cdots c_{\mathbf{r}_{K}}(t_{K})$ in slice $A$, we can express them as
@@ I.1 The case with fermion number conservation
I.2 Derivation for the case with pairing
Here we show how to diagonalize the non-negative Hermitian matrix $R$ defined in Eq. 36 in the Nambu basis $\Phi_{A}=(c_{\mathbf{r}_{1}}(t_{1}),\cdots,c_{\mathbf{r}_{K}}(t_{K}),c_{\mathbf{r}_{1}}^{\dagger}(t_{1}),\cdots,c_{\mathbf{r}_{K}}^{\dagger}(t_{K}))^{T}$:
$\{\Phi_{A},\Phi_{A}^{\dagger}\}=\left(\begin{array}[]{cc}B&V\\ V^{*}&B^{*}\end{array}\right)=R\ ,$ (S11)
where $B=B^{\dagger}$ is Hermitian, and $V=V^{T}$ is symmetric (not necessarily real). We mentioned in the main text that the system has a particle-hole symmetry $P=I_{K}\otimes\sigma_{x}$ due to the property of the Nambu basis:
$P\Phi_{A}=(\Phi_{A}^{\dagger})^{T}\ ,\qquad PRP^{-1}=R^{*}\ .$ (S12)
We now perform transformation to find a simplified basis. First, as we noted in the previous subsection, there exists a matrix $M$ satisfying $M^{\dagger}M=B^{-1}$, which indicates $MBM^{\dagger}=I$. So we have
$S_{1}RS_{1}^{\dagger}=\left(\begin{array}[]{cc}M&0\\ 0&M^{*}\end{array}\right)R\left(\begin{array}[]{cc}M^{\dagger}&0\\ 0&M^{T}\end{array}\right)=\left(\begin{array}[]{cc}I&MVM^{T}\\ M^{*}V^{*}M^{\dagger}&I\end{array}\right)\ .$ (S13)
Note that $MVM^{T}$ is still a symmetric matrix. Therefore, according to the Autonne-Takagi factorization, there exists a unitary matrix $U_{V}$ such that
$U_{V}(MVM^{T})U_{V}^{T}=\Lambda_{V}$ (S14)
is a real diagonal matrix with non-negative entries. We then define a matrix
$S_{2}=\left(\begin{array}[]{cc}U_{V}&0\\ 0&U_{V}^{*}\end{array}\right)\ ,$ (S15)
It is easy to see that
$S_{2}\left(\begin{array}[]{cc}I&MVM^{T}\\ M^{*}V^{*}M^{\dagger}&I\end{array}\right)S_{2}^{\dagger}=\left(\begin{array}[]{cc}I&\Lambda_{V}\\ \Lambda_{V}&I\end{array}\right)\ .$ (S16)
At this stage, if we define a new Nambu basis
$\Phi^{\prime}_{A}=S_{2}S_{1}\Phi_{A}=\left(\begin{array}[]{cc}U_{V}M&0\\ 0&U_{V}^{*}M^{*}\end{array}\right)\Phi_{A}\ ,$ (S17)
it will still satisfy the particle-hole symmetry $P\Phi^{\prime}_{A}=(\Phi_{A}^{\prime{\dagger}})^{T}$, since $PS_{2}P^{-1}=S_{2}^{*}$, $PS_{1}P^{-1}=S_{1}^{*}$. Therefore, we have
$\Phi^{\prime}_{A}=(f_{1},f_{2},\cdots,f_{K},f_{1}^{\dagger},f_{2}^{\dagger},\cdots,f_{K}^{\dagger})^{T}\ ,$ (S18)
where $f_{j}$ are fermionic linear operators (not yet fermion creation/annihilation operators) satisfying
$\{f_{i},f_{j}^{\dagger}\}=\delta_{ij}\ ,\qquad\{f_{i},f_{j}\}=\delta_{ij}\lambda_{j}\ ,$ (S19)
where we have assumed the real diagonal matrix $\Lambda_{V}$ has elements $\Lambda_{V}=\text{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{K})$, where $\lambda_{j}\geq 0$. Moreover, since
$2-2\lambda_{j}=2\{f_{j},f_{j}^{\dagger}\}-\{f_{j},f_{j}\}-\{f_{j}^{\dagger},f_{j}^{\dagger}\}=2A_{j}^{\dagger}A_{j}\geq 0\ ,\qquad\text{with }A_{j}=f_{j}-f_{j}^{\dagger}\ ,$ (S20)
(we have used the fact that any operator $A^{\dagger}A$ is non-negative) we know
$0\leq\lambda_{j}\leq 1\ ,\qquad\rightarrow\qquad 0\leq\Lambda_{V}\leq 1\ .$ (S21)
We can then define fermion opeartors $d_{j}$, $d_{j}^{\dagger}$ as follows:
$f_{j}=\sqrt{\frac{1+\sqrt{1-\lambda_{j}^{2}}}{2}}d_{j}+\sqrt{\frac{1-\sqrt{1-\lambda_{j}^{2}}}{2}}d_{j}^{\dagger}\ ,\qquad f_{j}^{\dagger}=\sqrt{\frac{1-\sqrt{1-\lambda_{j}^{2}}}{2}}d_{j}+\sqrt{\frac{1+\sqrt{1-\lambda_{j}^{2}}}{2}}d_{j}^{\dagger}\ ,$ (S22)
or more explicitly after solving the above equation:
$d_{j}=\sqrt{\frac{1+\sqrt{1-\lambda_{j}^{2}}}{2(1-\lambda_{j}^{2})}}f_{j}-\sqrt{\frac{1-\sqrt{1-\lambda_{j}^{2}}}{2(1-\lambda_{j}^{2})}}f_{j}^{\dagger}\ ,\qquad d_{j}^{\dagger}=\sqrt{\frac{1+\sqrt{1-\lambda_{j}^{2}}}{2(1-\lambda_{j}^{2})}}f_{j}^{\dagger}-\sqrt{\frac{1-\sqrt{1-\lambda_{j}^{2}}}{2(1-\lambda_{j}^{2})}}f_{j}\ ,$ (S23)
which satisfies
$\{d_{i},d_{j}^{\dagger}\}=\delta_{ij}\ ,\qquad\{d_{i},d_{j}\}=\{d_{i}^{\dagger},d_{j}^{\dagger}\}=0\ ,$ (S24)
namely, they are the usual orthogonal fermion annihilation and creation operators. More symbolically, we define the following matrix
$S_{3}=\left(\begin{array}[]{cc}s_{+}&-s_{-}\\ -s_{-}&s_{+}\end{array}\right)\ ,\qquad s_{\pm}=\sqrt{\frac{1\pm\sqrt{1-\Lambda_{V}^{2}}}{2(1-\Lambda_{V}^{2})}}\ .$ (S25)
We then have
$S_{3}S_{2}S_{1}RS_{1}^{\dagger}S_{2}^{\dagger}S_{3}^{\dagger}=LRL^{\dagger}=\left(\begin{array}[]{cc}I&0\\ 0&I\end{array}\right)\ ,\qquad L=S_{3}S_{2}S_{1}\ ,\qquad PLP^{-1}=L^{*}\ ,$ (S26)
namely, $R$ can be diagonalized by a particle-hole preserving transformation (note that $L$ is not unitary). We can rewrite the orthonormal Nambu basis
$\Psi_{A}=(d_{1},d_{2},\cdots,d_{K},d_{1}^{\dagger},d_{2}^{\dagger},\cdots,d_{K}^{\dagger})^{T}\ ,$ (S27)
which is related to the original basis via
$\Psi_{A}=L\Phi_{A}\ .$ (S28)
Note that Eq. S26 implies that the transformation matrix $L$ satisfies
$L^{\dagger}L=R^{-1}\ ,\qquad PLP^{-1}=L^{*}\ .$ (S29)